Hi, I don't know if this equation belongs here but I would like some help with this problem.
If f(x)=x^2 and g(x)=x+1, find f(4) + g(2)?
Any help would be appreciated! :)

Found 2 solutions by math_tutor2020, ikleyn:
Answer by math_tutor2020(3816)


Answer: 19

Work Shown:

f(x) = x^2
f(4) = 4^2
f(4) = 16

g(x) = x+1
g(2) = 2+1
g(2) = 3

f(4)+g(2) = 16+3
f(4)+g(2) = 19

Answer by ikleyn(52771)

.

Hello,

if you come to the forum with such questions, you are definitely in wrong class.